OpenStudy (anonymous):

Double check for me please? What is the 11th term in the sequence: an = -6n + 13? a. 79 b. -79 c. -53 d. 53 my answer: c. -53

OpenStudy (anonymous):

Hi again @ParthKohli is my answer right? haha

OpenStudy (shadowfiend):

Seems legit ;)

Parth (parthkohli):

n = 11 here. a11 = -6(11) + 13 a11 = -66 + 13 a11 = -53 CORRECTO!

OpenStudy (anonymous):

Awesome!! THank you!! ;) that's what i did and yay! it's right! :) woot woot!